# Low-Rank Online Dynamic Assortment with Dual Contextual Information

### v1

@@ Abstract
As e-commerce expands, delivering real-time personalized recommendations from vast catalogs poses a critical challenge for retail platforms. Maximizing revenue requires careful consideration of both individual customer characteristics and available item features to optimize assortments over time. In this paper, we consider the dynamic assortment problem with dual contexts – user and item features. In high-dimensional scenarios, the quadratic growth of dimensions complicates computation and estimation. To tackle this challenge, we introduce a new low-rank dynamic assortment model to transform this problem into a manageable scale. Then we propose an efficient algorithm that estimates the intrinsic subspaces and utilizes the upper confidence bound approach to address the exploration-exploitation trade-off in online decision making. Theoretically, we establish a regret bound of $\tilde{O}((d_{1}+d_{2})r\sqrt{T})$, where $d_{1},d_{2}$ represent the dimensions of the user and item features respectively, $r$ is the rank of the parameter matrix, and $T$ denotes the time horizon. This bound represents a substantial improvement over prior literature, made possible by leveraging the low-rank structure. Extensive simulations and an application to the Expedia hotel recommendation dataset further demonstrate the advantages of our proposed method.
@@ 1 Introduction
The primary objective of the platform is to maximize the total revenue over a time horizon $T$. It can adopt a ‘greedy strategy’ aimed at maximizing the revenue for each individual user by offering what is estimated as the best assortment at each time. However, due to the uncertainty of estimation, pure exploitation would lead to sub-optimal actions. To improve the accuracy of estimating these optimal assortments, the platform must delve into exploring user behaviors across various items. These two strategies, exploration for improving user preference estimation and exploitation for revenue maximization, often do not align in online decision making problems. Reinforcement learning, an area gaining increasing attention in statistics (Chakraborty and Murphy, , 2014; Zhao et al., , 2015; Shi et al., , 2018, 2023; Zhou et al., 2024b, ), provides valuable insights for navigating the delicate balance of the exploration-exploitation trade-off.
In recent years, a plethora of studies have utilized bandit and reinforcement learning algorithms (Lattimore and Szepesvári, , 2020) to address the dynamic assortment problem. Some research focuses on the non-contextual scenario (Chen and Wang, , 2018; Agrawal et al., , 2019), while others delve into the contextual setups (Chen et al., , 2020; Oh and Iyengar, , 2021; Goyal and Perivier, , 2022) under the multinomial logit (MNL) choice model (McFadden, , 1974) to characterize user behaviors. Advances in digital retail platform infrastructure have empowered these platforms with rich data on items and users, including demographics, search histories, and purchase records. By harnessing user features, the platform can tailor personalized assortments by looking back at the choice history of users with similar user contexts. Similarly, utilizing item features allows the platform to infer item preferences by analyzing the choice history of items sharing similar item contexts, even when the item has limited inclusion in assortment history. This motivates us to study the contextual dynamic assortment problem, specifically focusing on the incorporation of “dual contextual information” involving both items and users.
In contextual dynamic assortment, a common strategy for handling dual contexts is to stack the dual features as a joint feature vector and utilize its linear form (Sumida and Zhou, , 2023). Let $q_{t}\in\operatorname*{\mathbb{R}}^{d_{1}}$ be the vector representation of the feature of user $t$ and $p_{i}\in\operatorname*{\mathbb{R}}^{d_{2}}$ be a vector representation of the feature of item $i$. By stacking the dual features as $x_{it}=(1,p_{i}^{\top},q_{t}^{\top})^{\top}\in\operatorname*{\mathbb{R}}^{d_{1}+d_{2}+1}$, the utility of item $i$ for user $t$ is represented as $v_{it}=x_{it}^{\top}\phi$ with vector parameter $\phi\in\operatorname*{\mathbb{R}}^{d_{1}+d_{2}+1}$; see Figure 2.
@@
Another approach involves utilizing the vectorization of the combined effects of the two features as a joint feature vector, followed by the application of existing contextual dynamic assortment selection methods (Agrawal et al., , 2019; Oh and Iyengar, , 2021). By including the intercept and taking the outer product of the two extended features, it accounts for the intercept, main effects, and user-item interaction effects simultaneously; see Figure 3. However, in high-dimensional settings, the dimension of the outer product is $(d_{1}+1)\times(d_{2}+1)$, which scales with the product of the dimensions of the dual features, making the estimation of the effects computationally challenging.
@@
Inspired by this, in this paper we adopt the bilinear form $v=p^{\top}\Phi q$ with a low-rank matrix parameter $\Phi$ to represent utilities. This form offers flexibility, accommodating both stacking and vectorization approaches of incorporating dual features. The stack approach can be interpreted as the bilinear form with intercepts and a rank-2 parameter matrix while the vectorization approach can be interpreted as the scenario with full rank. Moreover, we can reduce the computational complexity by imposing a low-rank structure on $\Phi$, reducing the dimension of the parameter from $d_{1}d_{2}$ to $(d_{1}+d_{2})r$, where $r(\ll\min\{d_{1},d_{2}\})$ is the rank of the parameter matrix. Low-rank models have garnered growing interest in statistics (Jain et al., , 2013; Xia and Yuan, , 2021; Xia et al., , 2022; Ma et al., , 2024) with diverse scientific and business applications (Bell and Koren, , 2007; Koren et al., , 2009; Udell and Townsend, , 2019; Zhen and Wang, , 2024; Zhou et al., 2024a, ). Following this trend, our paper considers a new Low-rank Online Assortment with Dual-contexts (LOAD) problem.
@@
To solve the LOAD problem, we propose an Explore-Low-rank-Subspace-then-Apply-UCB (ELSA-UCB) algorithm which consists of two main stages: subspace exploration and the application of Upper Confidence Bound (UCB). The major computational bottleneck arises from estimating the entire preference parameter matrix. To overcome this challenge, we project the features and parameters to a low-dimensional subspace. In the first stage, we leverage the inherent low-rank structure to estimate the low-dimensional subspace of the matrix parameter. To accomplish this, we explore the space by employing random assortments and then solve the rank-constrained likelihood maximization problem to estimate the subspace. Given the non-convex nature of the rank-constrained optimization problem, we adopt the alternating gradient descent algorithm based on the Burer-Monteiro formulation (Burer and Monteiro, , 2003; Zheng and Lafferty, , 2016; Chi et al., , 2019) to estimate the parameter matrix. Subsequently, we perform singular value decomposition (SVD) on the estimated matrix to derive the intrinsic subspace. Utilizing the estimated subspace, we rotate the features and parameters in accordance with the subspace, then truncate negligible terms to effectively reduce the problem’s dimensionality. In the second stage, we employ the UCB-based strategy on the reduced dimension based on careful construction of the confidence bounds in this space. UCB uses an “optimism in the face of uncertainty” idea that addresses the exploration-exploitation trade-off by selecting the best action based on the “optimistic reward”, the upper confidence bound of the estimated reward (Auer, , 2002; Lattimore and Szepesvári, , 2020). Contrary to the unbiased model in the original parameter space, the preference model within the reduced space, with respect to the estimated subspaces is inherently biased. To handle this, we introduce a novel tool that provides a confidence bound for the expected reward on the reduced space, by correcting this bias. Implementing the UCB-based policy in this reduced space reduces the parameter dimension from $d_{1}\times d_{2}$ to $r(d_{1}+d_{2})-r^{2}$. This reduction markedly improves computational efficiency, reduces estimation error, and enhances the algorithm performance.
@@
To assess the theoretical performance of our proposed algorithm, we study the cumulative regrets of our policy. The regret is defined as the difference between the optimal reward from the oracle possessing complete knowledge of environmental parameters and the reward obtained from a given policy. In theory, we establish an $\tilde{O}((d_{1}+d_{2})r\sqrt{T})$ regret upper bound for our proposed ELSA-UCB policy (Theorem 1). Notably, in low-rank setups with $r\ll\min\{d_{1},d_{2}\}$, our regret bound is much improved from $\tilde{O}(d_{1}d_{2}\sqrt{T})$ regret bounds achieved by existing methods (Chen et al., , 2020; Oh and Iyengar, , 2021) that vectorize the outer product of the dual features.
@@
The regret analysis for our policy presents several challenges. Firstly, in the dual-contextual environment, the distribution of joint features depends on the assortment selection, violating the i.i.d. feature assumption in existing dynamic assortment selection literature (Chen et al., , 2020; Oh and Iyengar, , 2021). To accommodate this, we impose minimal assumptions on the item features (Assumption 2) to ensure the identifiability of parameters. Secondly, the non-convex nature of the low-rank optimization problem precludes us from leveraging classical convex optimization theories. While Jun et al., (2019); Kang et al., (2022) studied generalized linear bandits with low-rank structure, their analysis tools are not applicable to our LOAD problem due to the unique combinatorial structure in the dynamic assortment selection. Thus, we establish novel estimation bounds specifically tailored for the rank-constrained optimization problem on assortment scenarios.
@@
Finally, we present numerical results to demonstrate the effectiveness of our algorithm. Initially, we evaluate the ELSA-UCB policy’s performance on synthetic data, exploring various scenarios including changing ranks, feature dimensions, number of items, and assortment capacity. We compare cumulative regret over different time horizons with two variants of the UCB-MNL policy (Oh and Iyengar, , 2021): Stacked UCB-MNL and Vectorized UCB-MNL. Our analysis consistently shows that ELSA-UCB outperforms both methods across all scenarios, with the performance gap widening in scenarios with lower ranks and higher dimensions. Additionally, we conduct real data analysis on the Expedia dataset to optimize hotel recommendations. Our findings demonstrate that the parameter matrix exhibits a low-dimensional structure, and our method effectively leverages this to achieve improved performance, highlighting its practical usefulness.
@@ 1.1 Related Work
Our work is closely related to recent studies on contextual dynamic assortment selection and low-rank bandits. Additional relevant literature is provided in the Supplementary Materials.
1.1.1 Contextual Dynamic Assortment Selection
@@
Inspired by the foundational work by Caro and Gallien, (2007), numerous studies have explored the dynamic assortment problem under the MNL bandit model. Various policies, including Explore-Then-Commit (Rusmevichientong et al., , 2010; Sauré and Zeevi, , 2013), Thompson Sampling (Agrawal et al., , 2017) and UCB (Agrawal et al., , 2019) have been proposed to address the balance between exploration and exploitation in the MNL bandit model. Other studies consider various scenarios such as robustness against outliers (Chen et al., , 2023), multi-stage choices (Xu and Wang, , 2023) or extension to different choice models (Aouad et al., , 2023). Contextual dynamic assortment considers various facets, including item-specific features (Wang et al., , 2019), time-varying item features (Chen et al., , 2020), user types (Kallus and Udell, , 2020), item features (Shao et al., , 2022), or features associated with user-item pairs (Oh and Iyengar, , 2021; Goyal and Perivier, , 2022). These features are utilized in modeling utilities for the MNL bandit model. However, while existing algorithms address single-context scenarios, focusing solely on users or items, they lack the capability to estimate interactions between dual contexts. Although algorithms designed for joint features on user-item pairs could be applied to the dual contextual setup by interpreting the vectorization of the outer product of dual features, this approach inevitably encounters large computational burdens and fails to capture the low-rank structure of the parameter matrix in our LOAD problem. Our proposed method aims to fill this gap.
1.1.2 Low-Rank Bandits
@@
The dynamic environment involving high-dimensional covariates has also been a vibrant area of research. Based on a rich background of statistical tools for high dimensional problems (Wainwright, , 2019), numerous studies addressed the ‘curse of dimensionality’ in high-dimensional bandits. These studies often assume sparsity in the parameter vector and employ the LASSO-based approach (Abbasi-Yadkori et al., , 2012; Kim and Paik, , 2019) or presume the low-rank structure within the parameter matrix (Jun et al., , 2019; Kang et al., , 2022; Cai et al., , 2023; Zhou et al., 2024a, ). While certain studies have utilized the LASSO-based approach to address the sparse dynamic assortment problem (Wang et al., , 2019; Shao et al., , 2022), there has been comparatively less emphasis on leveraging the low-rank structure in the dynamic assortment. Recently, Cai et al., (2023) approached the dynamic assortment problem by modeling rewards as a bilinear form of action and context. However, they did not consider any user choice model and overlooked the inherent structure between assortment sets. Because of this, their algorithms and analysis tools are not applicable to our problem. To the best of our knowledge, our approach is the first one to incorporate the low-rank structure in the dynamic assortment selection with dual contexts.

### v2

@@ Abstract
As e-commerce expands, delivering real-time personalized recommendations from vast catalogs poses a critical challenge for retail platforms. Maximizing revenue requires careful consideration of both individual customer characteristics and available item features to continuously optimize assortments over time. In this paper, we consider the dynamic assortment problem with dual contexts – user and item features. In high-dimensional scenarios, the quadratic growth of dimensions complicates computation and estimation. To tackle this challenge, we introduce a new low-rank dynamic assortment model to transform this problem into a manageable scale. Then we propose an efficient algorithm that estimates the intrinsic subspaces and utilizes the upper confidence bound approach to address the exploration-exploitation trade-off in online decision making. Theoretically, we establish a regret bound of $\tilde{O}((d_{1}+d_{2})r\sqrt{T})$, where $d_{1},d_{2}$ represent the dimensions of the user and item features respectively, $r$ is the rank of the parameter matrix, and $T$ denotes the time horizon. This bound represents a substantial improvement over prior literature, achieved by leveraging the low-rank structure. Extensive simulations and an application to the Expedia hotel recommendation dataset further demonstrate the advantages of our proposed method.
@@ 1 Introduction
The primary objective of the platform is to maximize the total revenue over a time horizon $T$. It can adopt a ‘greedy strategy’ aimed at maximizing the revenue for each individual user by offering what is estimated as the best assortment at each time. However, due to the uncertainty of estimation, pure exploitation would lead to sub-optimal actions. To improve the accuracy of estimating these optimal assortments, the platform must delve into exploring user behaviors across various items. These two strategies, exploration for improving user preference estimation and exploitation for revenue maximization, often do not align in online decision making problems. Reinforcement learning, an area gaining increasing attention in statistics (Chakraborty and Murphy, 2014; Zhao et al., 2015; Shi et al., 2018, 2023; Zhou et al., 2024b), provides valuable insights for navigating the delicate balance of the exploration-exploitation trade-off.
In recent years, a plethora of studies have utilized bandit and reinforcement learning algorithms (Lattimore and Szepesvári, 2020) to address the dynamic assortment problem. Some research focuses on the non-contextual scenario (Chen and Wang, 2018; Agrawal et al., 2019), while others delve into the contextual setups (Chen et al., 2020; Oh and Iyengar, 2021; Goyal and Perivier, 2022) under the multinomial logit (MNL) choice model (McFadden, 1974) to characterize user behaviors. Advances in digital retail platform infrastructure have empowered these platforms with rich data on items and users, including demographics, search histories, and purchase records. By harnessing user features, the platform can tailor personalized assortments by looking back at the choice history of users with similar user contexts. Similarly, utilizing item features allows the platform to infer item preferences by analyzing the choice history of items sharing similar item contexts, even when the item has limited inclusion in assortment history. This motivates us to study the contextual dynamic assortment problem, specifically focusing on the incorporation of “dual contextual information” involving both items and users.
In contextual dynamic assortment, a common strategy for handling dual contexts is to stack the dual features as a joint feature vector and utilize its linear form (Sumida and Zhou, 2023). Let $q_{t}\in\operatorname*{\mathbb{R}}^{d_{1}}$ be the vector representation of the feature of user $t$ and $p_{i}\in\operatorname*{\mathbb{R}}^{d_{2}}$ be a vector representation of the feature of item $i$. By stacking the dual features as $x_{it}=(1,p_{i}^{\top},q_{t}^{\top})^{\top}\in\operatorname*{\mathbb{R}}^{d_{1}+d_{2}+1}$, the utility of item $i$ for user $t$ is represented as $v_{it}=x_{it}^{\top}\phi$ with vector parameter $\phi\in\operatorname*{\mathbb{R}}^{d_{1}+d_{2}+1}$; see Figure 2.
@@
Another approach involves utilizing the vectorization of the combined effects of the two features as a joint feature vector, followed by the application of existing contextual dynamic assortment selection methods (Agrawal et al., 2019; Oh and Iyengar, 2021). By including the intercept and taking the outer product of the two extended features, it accounts for the intercept, main effects, and user-item interaction effects simultaneously; see Figure 3. However, in high-dimensional settings, the dimension of the outer product is $(d_{1}+1)\times(d_{2}+1)$, which scales with the product of the dimensions of the dual features, making the estimation of the effects computationally challenging.
@@
Inspired by this observation, in this paper we adopt the bilinear form $v=p^{\top}\Phi q$ with a low-rank matrix parameter $\Phi$ to represent utilities. This form offers flexibility, accommodating both stacking and vectorization approaches of incorporating dual features. The stack approach can be interpreted as the bilinear form with intercepts and a rank-2 parameter matrix while the vectorization approach can be interpreted as the scenario with full rank. Moreover, we can reduce the computational complexity by imposing a low-rank structure on $\Phi$, reducing the dimension of the parameter from $d_{1}d_{2}$ to $(d_{1}+d_{2})r$, where $r(\ll\min\{d_{1},d_{2}\})$ is the rank of the parameter matrix. Low-rank models have garnered growing interest in statistics (Jain et al., 2013; Xia and Yuan, 2021; Xia et al., 2022; Ma et al., 2024) with diverse scientific and business applications (Bell and Koren, 2007; Koren et al., 2009; Udell and Townsend, 2019; Zhen and Wang, 2024; Zhou et al., 2024a). Following this trend, our paper considers a new Low-rank Online Assortment with Dual-contexts (LOAD) problem.
@@
To solve the LOAD problem, we propose an Explore-Low-rank-Subspace-then-Apply-UCB (ELSA-UCB) algorithm which consists of two main stages: subspace exploration and the application of Upper Confidence Bound (UCB). The major computational bottleneck arises from estimating the entire preference parameter matrix. To overcome this challenge, we project the features and parameters to a low-dimensional subspace. In the first stage, we leverage the inherent low-rank structure to estimate the low-dimensional subspace of the matrix parameter. To accomplish this, we explore the space by employing random assortments and then solve the rank-constrained likelihood maximization problem to estimate the subspace. Given the non-convex nature of the rank-constrained optimization problem, we adopt the alternating gradient descent algorithm based on the Burer-Monteiro formulation (Burer and Monteiro, 2003; Zheng and Lafferty, 2016; Chi et al., 2019) to estimate the parameter matrix. Subsequently, we perform singular value decomposition (SVD) on the estimated matrix to derive the intrinsic subspace. Utilizing the estimated subspace, we rotate the features and parameters in accordance with the subspace, then truncate negligible terms to effectively reduce the problem’s dimensionality. In the second stage, we employ the UCB-based strategy on the reduced dimension based on careful construction of the confidence bounds in this space. UCB uses an “optimism in the face of uncertainty” idea that addresses the exploration-exploitation trade-off by selecting the best action based on the “optimistic reward”, the upper confidence bound of the estimated reward (Auer, 2002; Lattimore and Szepesvári, 2020). Contrary to the unbiased model in the original parameter space, the preference model within the reduced space is inherently biased with respect to the estimated subspaces. To handle this, we introduce a novel tool that provides a confidence bound for the expected reward on the reduced space, by correcting this bias. Implementing the UCB-based policy in this reduced space reduces the parameter dimension from $d_{1}\times d_{2}$ to $r(d_{1}+d_{2})-r^{2}$. This reduction markedly improves computational efficiency, reduces estimation error, and enhances the algorithm performance. In practice, the rank of the underlying parameter matrix is often unknown. To address this issue, we propose ELSA-GIC, incorporating the rank selection procedure using Generalized Information Criterion (GIC) (Konishi and Kitagawa, 1996; Fan and Tang, 2013) into ELSA-UCB. We also prove rank selection consistency of ELSA-GIC in the LOAD problem.
@@
To assess the theoretical performance of our proposed algorithm, we study the cumulative regrets of our policy. The regret is defined as the difference between the optimal reward from the oracle possessing complete knowledge of environmental parameters and the reward obtained from a given policy. In theory, we establish an $\tilde{O}((d_{1}+d_{2})r\sqrt{T})$ regret upper bound for our proposed ELSA-UCB policy (Theorem 1). Notably, in low-rank setups with $r\ll\min\{d_{1},d_{2}\}$, our regret bound is much improved from $\tilde{O}(d_{1}d_{2}\sqrt{T})$ regret bounds achieved by existing methods (Chen et al., 2020; Oh and Iyengar, 2021) that vectorize the outer product of the dual features.
@@
The regret analysis for our policy presents several challenges. Firstly, in the dual-contextual environment, the distribution of joint features depends on the assortment selection, violating the i.i.d. feature assumption in existing dynamic assortment selection literature (Chen et al., 2020; Oh and Iyengar, 2021). To accommodate this, we impose minimal assumptions on the item features (Assumption 2) to ensure the identifiability of parameters. Secondly, the non-convex nature of the low-rank optimization problem precludes us from leveraging classical convex optimization theories. While Jun et al. (2019); Kang et al. (2022) studied generalized linear bandits with low-rank structure, their analysis tools are not applicable to our LOAD problem due to the unique combinatorial structure in the dynamic assortment selection. Thus, we establish novel estimation bounds specifically tailored for the rank-constrained optimization problem on assortment scenarios.
@@
Finally, we present numerical results to demonstrate the effectiveness of our algorithm. Initially, we evaluate the ELSA-GIC policy’s performance on synthetic data, exploring various scenarios including changing ranks, feature dimensions, number of items, and assortment capacity. We compare cumulative regret over different time horizons with two variants of the UCB-MNL policy (Oh and Iyengar, 2021): Stacked UCB-MNL and Vectorized UCB-MNL. Our analysis consistently shows that ELSA-GIC outperforms both methods across all scenarios, with the performance gap widening in scenarios with lower ranks and higher dimensions. We also show the robust performance of ELSA-GIC under model mis-specification. Additionally, we conduct real data analysis on the Expedia dataset to optimize hotel recommendations. Our findings demonstrate that the parameter matrix exhibits a low-dimensional structure, and our method effectively leverages this to achieve improved performance, highlighting its practical usefulness.
@@ 1.1 Related Work
Our work is closely related to recent studies on contextual dynamic assortment selection and low-rank bandits. Additional relevant literature is provided in Section S.1 of the supplement.
  • Contextual Dynamic Assortment Selection
@@
    Inspired by the foundational work by Caro and Gallien (2007), numerous studies have explored the dynamic assortment problem under the MNL bandit model. Various policies, including Explore-Then-Commit (Rusmevichientong et al., 2010; Sauré and Zeevi, 2013), Thompson Sampling (Agrawal et al., 2017) and UCB (Agrawal et al., 2019) have been proposed to address the balance between exploration and exploitation in the MNL bandit model. Other studies consider various scenarios such as robustness against outliers (Chen et al., 2023), multi-stage choices (Xu and Wang, 2023) or extension to different choice models (Aouad et al., 2023). Contextual dynamic assortment considers various facets, including item-specific features (Wang et al., 2019), time-varying item features (Chen et al., 2020), user types (Kallus and Udell, 2020), item features (Shao et al., 2022), or features associated with user-item pairs (Oh and Iyengar, 2021; Goyal and Perivier, 2022). These features are utilized in modeling utilities for the MNL bandit model. However, while existing algorithms address single-context scenarios, focusing solely on users or items, they lack the capability to estimate interactions between dual contexts. Although algorithms designed for joint features on user-item pairs could be applied to the dual contextual setup by interpreting the vectorization of the outer product of dual features, this approach inevitably encounters large computational burdens and fails to capture the low-rank structure of the parameter matrix in our LOAD problem. Our proposed method aims to fill this gap.
  • Low-Rank Bandits
@@
    The dynamic environment involving high-dimensional covariates has also been a vibrant area of research. Based on a rich background of statistical tools for high dimensional problems (Wainwright, 2019), numerous studies addressed the ‘curse of dimensionality’ in high-dimensional bandits. These studies often assume sparsity in the parameter vector and employ the LASSO-based approach (Abbasi-Yadkori et al., 2012; Kim and Paik, 2019) or presume the low-rank structure within the parameter matrix (Jun et al., 2019; Kang et al., 2022; Cai et al., 2023; Zhou et al., 2024a). While certain studies have utilized the LASSO-based approach to address the sparse dynamic assortment problem (Wang et al., 2019; Shao et al., 2022), there has been comparatively less emphasis on leveraging the low-rank structure in the dynamic assortment. Recently, Cai et al. (2023) approached the dynamic assortment problem by modeling rewards as a bilinear form of action and context. However, they did not consider any user choice model and overlooked the inherent structure between assortment sets. Because of this, their algorithms and analysis tools are not applicable to our problem. To the best of our knowledge, our approach is the first one to incorporate the low-rank structure in the dynamic assortment selection with dual contexts.
1.2 Notations
Let $[n]$ denote the set $\{1,2,\ldots,n\}$. The Frobenius norm of a matrix is denoted by $\|\cdot\|_{F}$, and the matrix-induced norm $\|v\|_{W}=\sqrt{v^{\top}Wv}$ is denoted by $\|\cdot\|_{W}$. The ball induced by the Frobenius norm, centered at $\Theta_{0}$ with radius $r$ is denoted by $\operatorname*{\mathcal{B}}(\Theta_{0},r):=\{\Theta:\|\Theta-\Theta_{0}\|_{F}\leq r\}$. For a vector $v$, $(v)_{i}$ represents its $i$-th component. We write $\Sigma_{1}\succ\Sigma_{2}$ if $\Sigma_{1}-\Sigma_{2}$ is positive definite. Denote the Kronecker product of two matrices by $\otimes$. We use the notation $\tilde{O}(a_{n})$ to denote that the quantity is bounded by $a_{n}$ up to logarithmic factors.